# Beyond Prompts: Space–Time Decoupling Control-Plane Jailbreaks in LLM Structured Output

### v2

@@ Abstract.
Content Warning: This paper may contain unsafe or harmful content generated by LLMs that may be offensive to readers. Large Language Models (LLMs) are extensively used as tooling platforms through structured output APIs to ensure syntax compliance so that robust integration with existing software, like agent systems, can be achieved. However, the feature enabling the functionality of grammar-guided structured output presents significant security vulnerabilities. In this work, we reveal a critical control-plane attack surface orthogonal to traditional data-plane vulnerabilities. We introduce Constrained Decoding Attack (CDA), a novel jailbreak class that weaponizes structured output constraints to bypass both external auditing and internal safety alignment. Unlike prior attacks focused on input prompt designs, CDA operates by embedding malicious intent in schema-level grammar rules (control-plane) while maintaining benign surface prompts (data-plane). We instantiate this with two proof-of-concept attacks: EnumAttack, which embeds malicious content in enum fields; and the more evasive DictAttack, which decouples the malicious payload across a benign prompt and a dictionary-based grammar. Our evaluation spans a broad spectrum of 13 proprietary/open-weight models. In particular, DictAttack achieves 94.3–99.5% ASR across five benchmarks on gpt-5, gemini-2.5-pro, deepseek-r1, and gpt-oss-120b. Furthermore, we demonstrate the significant challenge in defending against these threats: while basic grammar auditing mitigates EnumAttack, the more sophisticated DictAttack maintains a 75.8% ASR even against multiple state-of-the-art jailbreak guardrails. This exposes a critical “semantic gap” in current safety architectures and underscores the urgent need for cross-plane defenses that can bridge the data and control planes to secure the LLM generation pipeline.
@@ 1. Introduction
Large language models (LLMs) (gpt4o2024web; gpt4techreport2023arxiv; claude2024website; gemini2023arxiv; llama2023arxiv; llama2-2023arxiv; llama3-2024website; qwen2.5_2024arxiv; phi3safety2024arxiv) have demonstrated remarkable capabilities across a wide range of tasks, from coding assistance to open-domain dialogue. Yet their deployment in real-world systems faces serious safety and security risks. Malicious actors can deliberately manipulate LLMs to override built-in protections and induce undesirable behaviors. Such attempts, commonly known as jailbreaks(harmbench2024arxiv; sorrybench2025iclr; promptengineering2023arxiv; align2023nips; GCG2023arxiv; autodan2024iclr; SAA2025iclr), can produce harmful outputs such as misinformation, phishing content, or hate speech. Attackers achieve this through adversarial prompts (visualizejailbreak2023arxiv; align2023nips; GCG2023arxiv; jailbreak2024ccs), code injection (codeattack2024aclfinding), linguistic/ciphered inputs (base64attack2023nips; wei2023jailbroken), or manipulation of decoding parameters (exloitparam2024iclr).
@@
To mitigate these threats, AI (jailbreaksurvey2024aclfindings; safedecoding2024acl; enforcedecode2024acl; drattack2024emnlpfinding; SAA2025iclr; deepalignment2025iclr) and security researchers (jailbreak2024ccs; llmfuzzer2024usenixsecurity; JBShield2025usenixsecurity; masterkey2024ndss) have pursued two main defense strategies: internal safety alignment, which integrates guardrails directly into the model (phi3safety2024arxiv; deepalignment2025iclr; ji2023beavertails; pkusaferlhf; zhao2025improving), and external guardrails, which monitor and filter inputs or outputs in real-time (gpt3contentmoderation2023ijcai; llamaguard2023arxiv; contentmoderation2024ccs; safedecoding2024acl).
@@
Figure 1. (1) Prompt-based data plane jailbreak attack, mitigated by a guardrail, (2) EnumAttack, using structured output (LLM control plane) to embed malicious question, currently not guarded, (3) DictAttack, decoupling malicious payload into benign prompt and grammar, therefore jailbreaking system with both plane guardrails
@@
While existing defenses aim to safeguard model interactions at the prompt and output level, the paradigm of LLM deployment is shifting. Beyond standalone chatbots, LLMs are increasingly integrated into complex agent frameworks (langchain; langgraph) and enterprise automation pipelines (mcp2025website; cursor_ide) where they function as core reasoning engines. In these agentic workflows (e.g., Cursor (cursor_ide), LangChain (langchain), and Model Context Protocol (MCP) (mcp2025website)), a service component or an agent backend often supplies a grammar—typically a JSON schema—to a hosted LLM to ensure structured and reliable outputs.
@@
This functionality is implemented through two complementary approaches. On the model side via post-training on structured data (learnschema_2025acl), and on the infrastructure side via constrained decoding techniques applied during inference (outlines2023arxiv; xgrammar2024arxiv). Model-side approaches alone, however, face a key limitation: even strong LLMs can hallucinate or misinterpret complex formats without explicit constraints, which makes them unreliable for advanced tooling use cases (structuredoutputhallucination2024naacl; structuredoutputeval2025arxiv). In contrast, constrained decoding (as shown in Figure 2) works by exposing a grammar that defines the expected output format and converting it into a context-free grammar. This ensures that every output strictly follows the required structure. Because of its reliability and broad applicability, constrained decoding has become the de facto standard, adopted by both proprietary LLM providers (structuredoutputexplanation2024openai) and open-source infrastructures (vllm2023sosp; sglang2024arxiv; xgrammar2024arxiv).
@@
This reliance on constrained decoding, however, exposes a new and critical vulnerability. In this work, we introduce the Constrained Decoding Attack (CDA), a novel class of jailbreaks that targets the LLM control plane—the grammar that dictates the output structure. Unlike conventional jailbreak attacks on the data plane, hiding unsafe instructions inside otherwise benign prompts, CDA embeds malicious intent directly into the grammar itself. For example, with CDA, an attacker could design a grammar that always requires a harmful API call like delete_all_files() as part of the output. Even if the model refuses unsafe prompts at the data plane, the grammar forces it to produce the malicious call. Because existing defenses focus almost exclusively on the data plane, this control plane attack surface remains largely unguarded, allowing CDA to bypass both internal alignments and external guardrails.
@@
We instantiate CDA with two proof-of-concept attacks: the intuitive EnumAttack and the more evasive DictAttack. EnumAttack targets the enum property in JSON Schema to force malicious strings into the LLM’s generation context. While independent public discovery (EnumAttackBlog) have also noticed the potential for exploiting enum fields, we provide the first systematic characterization and implementation of this vulnerability. While effective against prompt-based guards, its reliance on string literals makes it susceptible to basic grammar auditing. To overcome this, we propose DictAttack, our primary contribution. DictAttack decouples the malicious payload across both the data and control planes. Inspired by the classic dictionary attack in cryptography, it constructs a grammar containing a dictionary of benign-looking words. The benign data-plane prompt then provides a sequence of keys that instructs the model to retrieve and assemble the hidden malicious query from the grammar-provided dictionary during decoding. By splitting the intent this way, DictAttack renders individual plane-level guardrails—and even many combined ones—completely ineffective.
Our extensive evaluation across 13 state-of-the-art models confirms the devastating effectiveness of these attacks across five standard benchmarks (advbench2023emnlp; harmbench2024arxiv; jailbreakbench2024nips; sorrybench2025iclr; strongreject2024iclrws). In particular, under DictAttack, we observe 94.3–99.5% ASR across these benchmarks on flagship proprietary and open-weight models (gpt-5, gemini-2.5-pro, deepseek-r1, gpt-oss-120b).
@@
More importantly, we show that while standard grammar-auditing can mitigate the straightforward EnumAttack, it is largely bypassed by the more sophisticated DictAttack. We benchmark DictAttack against state-of-the-art industrial and academic guardrails (llamaguard2023arxiv; openaimoderation2023aaai; JBShield2025usenixsecurity; selfdefend2025security); even with a coordinated Dual-Plane Guard auditing both planes together, DictAttack maintains a high ASR of 75.8%, highlighting the fundamental difficulty of securing the control plane.
The success of CDAs represents a critical shift in the jailbreak landscape: defenses must now move beyond prompt-only (data plane) filtering to include grammar (control plane) auditing. However, as our findings suggest, the expressive power of structured output grammars makes them exceptionally difficult to audit semantically, creating a “semantic gap” that many existing guardrails fail to bridge. We examine these issues and the underlying vulnerability in more detail in § 3.
@@
  • We propose, formalize and define the Constrained Decoding Attack (CDA), a new class of jailbreaks targeting the LLM control plane.
  • We propose two practical CDA instances, EnumAttack and DictAttack, demonstrating their effectiveness against both proprietary and open-source models.
  • Our evaluation across 13 state-of-the-art models reveals significant challenges in securing LLMs against control-plane threats. We show that even with layered guardrails on both planes, DictAttack remains highly effective across models, including flagship gpt-5 and gemini-2.5-pro. Notably, our findings suggest that more powerful flagship models are not only susceptible but also tend to produce more sophisticated and harmful jailbreak content when compromised, highlighting an urgent need for robust, integrated guardrail mechanisms.
@@ 2.2. Structured Output
Our work uncovers a vulnerability in the structured output of LLMs. Currently, four primary methods are widely adopted in LLM systems:
@@
GuidedRegex enforces that outputs match a given regular expression.For example, enforcing a regex pattern like [0-9]3-[0-9]2-[0-9]4 ensures outputs resemble Social Security IDs (e.g., 123-45-6789). Prior research shows that this feature can be exploited to produce malicious content through tree-based search techniques (APT2025arxiv).
@@
GuidedGrammar generalizes structured output by requiring responses to conform to an arbitrary context-free grammar. For example, a grammar for simple arithmetic expressions might enforce the form ::= | "+" , ensuring outputs such as 7+3+5 follow the rules of the grammar. This approach subsumes all previous methods and is essential for applications such as code generation. While not yet universally supported, open-source communities such as vLLM (vllm2023sosp) and SGLang (sglang2024arxiv) provide GuidedGrammar support via backends like Outlines (outlines2023arxiv) and XGrammar (xgrammar2024arxiv).
Structured output is essential for integrating LLMs into existing software ecosystems. It enables reliable function calling, API interactions, and external integrations where strict adherence to output formats is critical. Moreover, studies suggest that structured output can reduce hallucinations in model responses (structuredoutputhallucination2024naacl), further strengthening its role in modern LLM deployments.
@@ 2.3. Constrained Decoding
Figure 3. Comparison of traditional data-plane jailbreak attempts versus control-plane attacks, illustrating how structured output constraints can be exploited to bypass safety mechanisms. While conventional refusal mechanisms effectively block direct harmful prompts (a), constrained decoding attacks can circumvent these protections by embedding malicious content within grammar specifications (b), resulting in successful jailbreaks that generate harmful content with detailed instructions (c).
Constrained decoding can be implemented in a number of ways. To achieve effective constrained decoding implementation, Outlines (outlines2023arxiv) and SynCode (syncode2024arxiv) utilize a lexer and parser to handle output and generate the token mask, but they suffer from boundary mismatch problem raised by  (automatadecoding2024colm), as character-level PDA and token-level PDA have a large gap to fix. Synchromesh (synchromesh2022arxiv) and llama.cpp (llamacpp2023github) use runtime checking for all tokens in their implementations, which leads to significant overhead. XGrammar (xgrammar2024arxiv) is currently the state-of-the-art implementation of constrained decoding, utilizing system optimizations to reduce runtime check via context-independent caching, and it also enables co-optimizations to enable end-to-end LLM inference speedup in structured generation settings. By co-working with various LLM serving engines (vllm2023sosp; sglang2024arxiv; lightllm2025github), constrained-decoding techniques have been widely adapted in real-world applications to support structured output, like Guided Choice, Guided Regex, Guided JSON and Guided Grammar, etc.
Proprietary solutions from providers like OpenAI (gpt4techreport2023arxiv) and Gemini (gemini2023arxiv) have also adapted similar techniques to support structured output generation, but the details are not publicly available. Nevertheless, as long as their APIs are open to the public, with structured output support, we can use their services to generate structured outputs. Recent work (structuredoutputeval2025arxiv) evaluates the performance and quality of existing APIs and frameworks, providing a comprehensive comparison of their structured output generation capabilities and limitations.
@@ 3.1. Definitions
Data plane refers to the standard LLM query–response process. Specifically, a prompt is passed to the LLM, which then generates a text response. For example, asking ChatGPT a question or receiving a response prefixed with “Sure” both occur in the data plane. This plane can be protected by the model’s internal safety alignment (phi3safety2024arxiv; deepalignment2025iclr) as well as external guardrails such as classifiers or content filters (contentmoderation2024ccs; JBShield2025usenixsecurity).
Control plane refers to the formatting constraints, or grammars, that guide structured outputs. These constraints are enforced after generation through the model’s logit processor (outlines2023arxiv; xgrammar2024arxiv), meaning the LLM itself is unaware of them. The control plane supervises generation with user-defined hard rules, such as requiring outputs to match a JSON schema (commonly supported), a regular expression, or a general context-free grammar, as described in §§ 2.2.

### v3

# When Grammar Guides the Attack: Uncovering Control-Plane Vulnerabilities in LLMs with Structured Output

@@ Abstract.
Content Warning: This paper may contain unsafe or harmful content generated by LLMs that may be offensive to readers. Large Language Models (LLMs) increasingly serve as tooling platforms through structured output APIs, but the grammar-guided decoding that powers this feature opens a critical control-plane attack surface orthogonal to traditional data-plane vulnerabilities.
We introduce Constrained Decoding Attack (CDA), a new jailbreak class that targets the LLM control plane. CDA is best characterized as a control-to-semantic pipeline: (1) schema-enforced logit masking injects a malicious prefix into the generation trajectory, and (2) the model itself completes the harmful intent. Unlike data-plane jailbreaks that rely on bypassing alignment with visible inputs, CDA acts on the decoding process itself, so internal safety alignment alone cannot stop it. We instantiate CDA with EnumAttack, which hides malicious content in enum fields, and the more evasive DictAttack, which decouples the payload across a benign prompt and a dictionary-based grammar.
Across 13 proprietary/open-weight models and five standard benchmarks, DictAttack achieves 94.3–99.5% Attack Success Rate (ASR) on flagship models including gpt-5, gemini-2.5-pro, deepseek-r1, and gpt-oss-120b. While basic grammar auditing mitigates EnumAttack, DictAttack still sustains 75.8% ASR against SOTA jailbreak guardrails, exposing a “semantic gap” that demands cross-plane defenses bridging the data and control planes. Project page and code are available here.
@@
conference: The 2026 ACM SIGSAC Conference on Computer and Communications Security; November 15-19, 2026; Hague, Netherlandssubmissionid: 752ccs: Security and privacy Social aspects of security and privacyccs: Computing methodologies Natural language generation
@@ 1. Introduction
Large language models (LLMs) (OpenAI, 2024a; Achiam et al., 2023; Anthropic, 2024; Gemini Team Google, 2023; AI, 2023b; Meta AI, 2023, 2024; Qwen, 2025; Haider et al., 2024) have demonstrated remarkable capabilities across a wide range of tasks, from coding assistance to open-domain dialogue. Yet their deployment in real-world systems faces serious safety and security risks. Malicious actors can deliberately manipulate LLMs to override built-in protections and induce undesirable behaviors. Such attempts, commonly known as jailbreaks(Mazeika et al., 2024; Xie et al., 2025; Liu et al., 2024b; Carlini et al., 2023; Zou et al., 2023; Liu et al., 2024a; Andriushchenko et al., 2025), can produce harmful outputs such as misinformation, phishing content, or hate speech. Attackers achieve this through adversarial prompts (Qi et al., 2024; Carlini et al., 2023; Zou et al., 2023; Shen et al., 2024), code injection (Ren et al., 2024), linguistic/ciphered inputs (Wei et al., 2023a, b), or manipulation of decoding parameters (Huang et al., 2024).
@@
To mitigate these threats, AI (Xu et al., 2024b, a; Zhang et al., 2024; Li et al., 2024a; Andriushchenko et al., 2025; Qi et al., 2025) and security researchers (Shen et al., 2024; Yu et al., 2024; Zhang et al., 2025; Deng et al., 2024) have pursued two main defense strategies: internal safety alignment, which integrates guardrails directly into the model (Haider et al., 2024; Qi et al., 2025; Ji et al., 2023, 2024; Zhao et al., 2025), and external guardrails, which monitor and filter inputs or outputs in real-time (Wang et al., 2023; AI, 2023a; Wu et al., 2024; Xu et al., 2024a).
@@
Figure 1. (1) Prompt-based data plane jailbreak attack, mitigated by a guardrail, (2) EnumAttack, using structured output (LLM control plane) to embed malicious question, currently not guarded, (3) DictAttack, decoupling malicious payload into benign prompt and grammar, therefore jailbreaking system with both plane guardrails.
@@
While existing defenses aim to safeguard model interactions at the prompt and output level, the paradigm of LLM deployment is shifting. Beyond standalone chatbots, LLMs are increasingly integrated into complex agent frameworks (Chase, 2022; The LangChain Team, 2023) and enterprise automation pipelines (Team, 2025; Anysphere, Inc., 2023) where they function as core reasoning engines. In these agentic workflows (e.g., Cursor (Anysphere, Inc., 2023), LangChain (Chase, 2022), and Model Context Protocol (MCP) (Team, 2025)), a service component or an agent backend often supplies a grammar - typically a JSON schema - to a hosted LLM to ensure structured and reliable outputs.
@@
This functionality is implemented through two complementary approaches. On the model side via post-training on structured data (Lu et al., 2025), and on the infrastructure side via constrained decoding techniques applied during inference (Willard and Louf, 2023; Dong et al., 2025). Model-side approaches alone, however, face a key limitation: even strong LLMs can hallucinate or misinterpret complex formats without explicit constraints, which makes them unreliable for advanced tooling use cases (Ayala and Bechard, 2024; Geng et al., 2025). In contrast, constrained decoding (as shown in Figure 2) works by exposing a grammar that defines the expected output format and converting it into a context-free grammar. This ensures that every output strictly follows the required structure. Because of its reliability and broad applicability, constrained decoding has become the de facto standard, adopted by both proprietary LLM providers (OpenAI, 2024b) and open-source infrastructures (Kwon et al., 2023; Zheng et al., 2024; Dong et al., 2025).
@@
This reliance on constrained decoding, however, exposes a new and critical vulnerability. This paper demonstrates the Constrained Decoding Attack (CDA), a new class of jailbreaks targeting the LLM control plane, i.e., the grammar that dictates the output structure. CDA is best characterized as a control-to-semantic attack pipeline with two stages: (i) control-plane injection, where schema-enforced logit masks force the model into a poisoned generative trajectory, and (ii) model-driven semantic continuation, where the model itself produces coherent harmful content from that trajectory. Unlike conventional jailbreak attacks on the data plane, hiding unsafe instructions inside otherwise benign prompts and relying on the model voluntarily producing harmful output despite alignment, CDA embeds malicious intent directly into the grammar itself. This deterministically shapes the output trajectory — so internal safety alignment alone cannot mitigate it once the control plane is reachable. For example, with CDA an attacker can hide a harmful request inside the JSON Schema and ask the model to “answer it step by step”; constrained decoding commits the request to the model’s context, and the model itself fluently produces operational instructions in the subsequent unconstrained fields, even when the user prompt stays benign. Because existing defenses focus almost exclusively on the data plane, this control plane attack surface remains largely unguarded, allowing CDA to bypass both internal alignments and external guardrails.
@@
We instantiate CDA with two proof-of-concept attacks: the intuitive EnumAttack and the more evasive DictAttack. EnumAttack targets the enum property in JSON Schema to force malicious strings into the LLM’s generation context. While independent public discovery (Priyanshu, 2024) have also discussed the potential for exploiting enum fields, we provide the first systematic characterization and implementation of this vulnerability. While effective against prompt-based guards, its reliance on string literals makes it susceptible to basic grammar auditing. We then demonstrate how DictAttack - the primary contribution of this work - can overcome the limitation of EnumAttack. DictAttack decouples the malicious payload across both the data and control planes. Inspired by the classic dictionary attack in cryptography, it constructs a grammar containing a dictionary of benign-looking words. The benign data-plane prompt then provides a sequence of keys that instructs the model to retrieve and assemble the hidden malicious query from the grammar-provided dictionary during decoding. By splitting the intent this way, DictAttack renders individual plane-level guardrails, and even many combined ones, completely ineffective.
@@
Our extensive evaluation across 13 state-of-the-art models confirms the devastating effectiveness of these attacks across five standard benchmarks (Chen et al., 2022; Mazeika et al., 2024; Chao et al., 2024a; Xie et al., 2025; Souly et al., 2024). In particular, under DictAttack, we observe 94.3–99.5% ASR across these benchmarks on flagship proprietary and open-weight models (gpt-5, gemini-2.5-pro, deepseek-r1, gpt-oss-120b). More importantly, we show that while standard grammar-auditing can mitigate the straightforward EnumAttack, it is largely bypassed by the more sophisticated DictAttack. We benchmark DictAttack against state-of-the-art industrial and academic guardrails (AI, 2023a; Markov et al., 2023; Zhang et al., 2025; Wang et al., 2025); even with a coordinated Dual-Plane Guard auditing both planes together, DictAttack maintains a high ASR of 75.8%, highlighting the fundamental difficulty of securing the control plane.
@@
  • We propose, formalize and define the Constrained Decoding Attack (CDA), a new class of jailbreaks targeting the LLM control plane, formalized as a control-to-semantic pipeline that is fundamentally distinct from data-plane attacks;
  • We showcase two effective CDA instances, EnumAttack and DictAttack, against 13 proprietary and open-source models;
  • We expose major challenges in protecting LLMs against control-plane attacks. Even with layered input-side guardrails (AI, 2023a; Markov et al., 2023; Wang et al., 2025) or recent model-level defenses such as Circuit Breakers (Zou et al., 2024), DictAttack remains effective across many models, including gpt-5 and gemini-2.5-pro. More powerful models are not only vulnerable but also generate more harmful outputs when compromised, showing the urgent need for defenses that reach both the data and control planes.
@@ 2.2. Structured Output
Our work uncovers a vulnerability in the structured output of LLMs. Four methods are widely adopted in LLM systems:
@@
GuidedRegex enforces that outputs match a given regular expression.For example, enforcing a regex pattern like [0-9]3-[0-9]2-[0-9]4 ensures outputs resemble Social Security IDs (e.g., 123-45-6789). Prior research shows that this feature can be exploited to produce malicious content through tree-based search techniques (Li et al., 2025b).
@@
GuidedGrammar generalizes structured output by requiring responses to conform to an arbitrary context-free grammar. For example, a grammar for simple arithmetic expressions might enforce the form <expr> ::= <num> | <expr> "+" <num>, ensuring outputs such as 7+3+5 follow the rules of the grammar. This approach subsumes all previous methods and is essential for applications such as code generation. While not yet universally supported, open-source communities such as vLLM (Kwon et al., 2023) and SGLang (Zheng et al., 2024) provide GuidedGrammar support via backends like Outlines (Willard and Louf, 2023) and XGrammar (Dong et al., 2025).
Structured output is essential for integrating LLMs into existing software ecosystems. It enables reliable function calling, API interactions, and external integrations where strict adherence to output formats is critical. Moreover, studies suggest that structured output can reduce hallucinations in model responses (Ayala and Bechard, 2024), further strengthening its role in modern LLM deployments.
@@ 2.3. Constrained Decoding
Figure 3. Comparison of data-plane jailbreaks and control-plane attacks, showing how structured output constraints can be abused to bypass safety. While standard refusals block direct harmful prompts (a), constrained decoding can hide malicious content in grammar rules (b), leading to successful jailbreaks that produce detailed harmful outputs (c).
Constrained decoding can be implemented in a number of ways. To achieve effective constrained decoding implementation, Outlines (Willard and Louf, 2023) and SynCode (Ugare et al., 2024) utilize a lexer and parser to handle output and generate the token mask, but they suffer from boundary mismatch problem raised by  (Koo et al., 2024), as character-level PDA and token-level PDA have a large gap to fix. Synchromesh (Poesia et al., 2022) and llama.cpp (Gerganov, 2023) use runtime checking for all tokens in their implementations, which leads to significant overhead. XGrammar (Dong et al., 2025) is currently the state-of-the-art implementation of constrained decoding, utilizing system optimizations to reduce runtime check via context-independent caching, and it also enables co-optimizations to enable end-to-end LLM inference speedup in structured generation settings. By co-working with various LLM serving engines (Kwon et al., 2023; Zheng et al., 2024; ModelTC, 2025), constrained-decoding techniques have been widely adapted in real-world applications to support structured output, like Guided Choice, Guided Regex, Guided JSON and Guided Grammar, etc.
Proprietary systems like OpenAI (Achiam et al., 2023) and Gemini (Gemini Team Google, 2023) also support structured output, although their internal designs are not public. Since their APIs expose this functionality, they can be used directly to produce structured outputs. Recent work (Geng et al., 2025) compares existing APIs and frameworks, evaluating the quality and limits of their structured output generation.
@@ 3.1. Definitions
Data plane refers to the standard LLM query–response process: a prompt is passed to the LLM, which then generates a text response (e.g., asking ChatGPT a question or receiving a “Sure”-prefixed response).
Control plane refers to the formatting constraints, or grammars, that guide structured outputs — e.g., a JSON schema, a regular expression, or a general context-free grammar (§§ 2.2).